# Matteo Escudé

### v5

@@ 3 Equilibrium skepticism and unraveling
Being able to prove news better than the prior means that the sender can—by acquiring information in a fashion that makes sending $m_{h}$ possible with positive probability—induce in the receiver a belief that she strictly prefers over the prior. Importantly, this is possible irrespective of the receiver’s belief (i.e. even if she is maximally skeptical, assigning belief $\min M^{-1}(m_{h})$ to message $m_{h}$).
@@
Observe that PNBP can hold only if, in the terminology of Kamenica and Gentzkow (2011), there is information the sender would share i.e. only if the sender is not already obtaining her highest possible payoff at the prior. [^24] That is: only if $v(\overline{p})<v(1)$, since $v$ is non-decreasing. Notice that in full-verifiability settings (Examples 1 and 2) this condition is also sufficient for PNBP. In contrast, in cheap-talk environments (Example 3) the sender cannot prove news better than the prior for any admissible $v$ and $\overline{p}$ (so even if there is information she would share).
Equipped with these definitions, we can state the first result, which describes the properties of every equilibrium when the sender can prove news better than the prior. The proof is in Appendix B.
Theorem 1.
If the sender can prove news better than the prior, in any equilibrium every on-path sender type is lowest-consistent with any message it sends with positive probability.
The sender therefore never acquires better news than what she only just needs to send each of the messages used in equilibrium. The result also implies that every message sent with positive probability must come from the lowest type able to send that message. It follows that only types that are lowest-consistent with some message can be on the equilibrium path.
Theorem 1 implies the following corollary, which sharply characterizes equilibrium behavior and constitutes the central result of this section. To state it, I first formally define ‘maximal skepticism’ and ’full revelation of information’ in the context of the model.
Definition 4.
Given a sender strategy $(\pi,\mu)$ and receiver belief $\beta$, say that the receiver is maximally skeptical following message $m$ iff $\beta(m)=\min M^{-1}(m)$. We say that the sender reveals all acquired information iff for every $s\in\operatorname*{supp}\pi_{S}$ and every $m\in\operatorname*{supp}\mu(m|\pi,s)$, $\beta(m)=s$.
Corollary 1 (Skepticism & full revelation).
If the sender can prove news better than the prior, in any equilibrium:
  1. (1)
    the receiver is maximally skeptical following any on-path message,
  2. (2)
    the sender reveals all acquired information.
Proof.
For point (1), consider any on-path message $m$. Theorem 1 implies that $m$ is sent by type $\min M^{-1}(m)$ only. Beliefs are to satisfy Bayes’ rule on-path, so $\beta(m)=\min M^{-1}(m)$. For point (2), consider any on-path type $s$. Theorem 1 implies that, for any message $m$ sent by type $s$ with positive probability, it must be that $s=\min M^{-1}(m)$. Since $\beta(m)=\min M^{-1}(m)$ from (1) above, we have $\beta(m)=s$ for every message $m$ sent with positive probability by any on-path type $s$, as required.
The intuition behind the result can be summarized as follows. If the receiver were not maximally skeptical following an on-path message, the sender could profitably deviate to a signal ‘worsening the meaning’ of that message. Such a deviation can always be constructed, unless the sender only acquires information by assigning positive probability to lowest-consistent types.
This logic is related to the force driving the results in Ben-Porath et al. (2018) and DeMarzo et al. (2019). In these papers, however, the only non-verifiable message whose meaning must be determined in equilibrium is ‘non-disclosure’.  [^25] They work in Dye (1985)-like environments, which prevents the equilibrium from fully unraveling. This leads the sender to take an equilibrium action which spreads the distribution of outcomes (in Ben-Porath et al. (2018)) or which minimizes her payoff following non-disclosure (in DeMarzo et al. (2019)). In contrast, in my model, the meaning of any message $m$ for which $M^{-1}(m)$ is not a singleton is determined in equilibrium.
As discussed in the introduction, the model combines two modes of persuasion: selective information production and misrepresentation of the findings. Corollary 1 demonstrates that—under a mild condition on what the sender can prove—only the former persuasion channel is active in equilibrium, as the receiver fully learns what the sender discovered. How the sender can misreport, however, shapes what the sender chooses to learn, as she will only acquire information in ways that allow her to credibly separate from all lower types in equilibrium. Thus, while muted in equilibrium, the ability to misreport directly influences how the sender chooses to acquire information.
Observe also that in evidence games with partial verifiability (for example, Okuno-Fujiwara et al. (1990) and Hagenbach et al. (2014)) it is not the case that the sender reveals all of her private information: there are type distributions such that distinct sender types pool together by sending the same messages, obtaining the same payoff in equilibrium. In my model—as a consequence of the sender’s type distribution being an equilibrium object—this cannot occur. Put differently, for any type distribution such that pooling would occur in the communication stage, there is a profitable deviation for the sender in the information acquisition stage.
The following example further illustrates the logic behind Theorem 1 and Corollary 1 and how, despite the sender revealing all acquired information, the receiver still only obtains partial information about the state.
3.1 A three-action example
Let $v$ be defined as follows:
$v(s)=\begin{cases}0\quad\text{if}\quad s<2/5,\\ 1\quad\text{if}\quad s\in[2/5,4/5),\\ 3\quad\text{if}\quad s\geq 4/5.\end{cases}$
Fix the prior at $\overline{p}=1/3$. Consider the following ‘certifiable thresholds’ verifiability structure (as introduced in Example 5) with just two messages: $m_{L}$ and $m_{M}$. $m_{L}\in M(s)$ for each $s\in[0,1]$ (since it is available to all types, this can be interpreted as a ‘silent’ or ‘I don’t know’ message) while $m_{M}\in M(s)$ iff $s\in[1/2,1]$ i.e. $m_{M}$ proves to the receiver that the state has at least a 50% chance of being $1$.
Since an equilibrium exists (this is proven for the general case in Appendix A) and since the sender can prove news better than the prior (with message $m_{M}$), Theorem 1 implies that in the only equilibrium the sender chooses a signal ‘splitting’ the prior into posteriors 00 and $1/2$ (the only two lowest-consistent types). The theorem further implies that following realization 00, she sends message $m_{L}$, and following realization $1/2$, she sends $m_{M}$. Observe finally, as highlighted in Corollary 1, that the sender reveals all acquired information—the receiver’s on-path posterior beliefs coincide with the sender’s ($\beta(m_{L})=0$ and $\beta(m_{M})=1/2$)—but the value of the state is not perfectly revealed to the receiver (as there is uncertainty following message $m_{M}$).
[…]
Figure 1: Sender payoff function $v$ (bold line) and verifiability structure (below the horizontal axis) in the three-action example.
Why can’t a signal assigning positive probability only to $\{0,4/5\}$, for example, be an equilibrium? Note that this would be the equilibrium signal if the sender’s information acquisition were public, as in Kamenica and Gentzkow (2011). [^26] In the Kamenica and Gentzkow (2011) case, the sender’s equilibrium payoff would be $(\operatorname*{cav}v)(1/3)$, which she would attain by choosing a signal splitting the prior between 00 and $4/5$, as can be deduced by inspecting Figure 1. Suppose that, indeed, the sender chose such a signal in equilibrium, sending message $m_{L}$ following 00 and $m_{M}$ following $4/5$. In the candidate equilibrium, the sender would obtain a payoff of 00 following message $m_{L}$, and a payoff of $3$ following message $m_{M}$. However, the sender could deviate to a signal that splits the prior between 00 and $4/5-\delta$ (for $\delta<10/3$); in the former case sending message $m_{L}$ and in the latter sending message $m_{M}$. This deviation is not detectable by the receiver and it increases (decreases) the probability of sending $m_{M}$ ($m_{L}$), thus strictly increasing the sender’s expected payoff. That is: since $4/5$ is not lowest-consistent, the sender can profitably ‘worsen the meaning’ of $m_{M}$, without being detected.
3.2 Value of persuasion
Theorem 1 also directly implies that the value of optimal persuasion in this context is unique, and that it can be characterized as the concave envelope of a skepticism-adjusted version of $v$. Letting
$v_{-}(s)\equiv v\left(\max_{m\in M(s)}\min M^{-1}(m)\right),$ (1)
for each $s\in[0,1]$, we have the following further corollary of Theorem 1.
Corollary 2 (Unique value of persuasion).
If the sender can prove news better than the prior, in any equilibrium the sender’s expected payoff is $(\operatorname*{cav}v_{-})(\overline{p})$.
That is, the sender’s equilibrium expected payoff is the concave envelope of the ‘value of interim sender beliefs’, when facing a receiver who is maximally skeptical following any message.
Proof.
Consider any equilibrium $(\pi,\mu,\beta)$. Let $w_{\beta}(s)\equiv\max_{m\in M(s)}v(\beta(m))$, $s\in[0,1]$. The sender’s equilibrium expected payoff is $(\operatorname*{cav}w_{\beta})(\overline{p})$, since—for a given $\beta$ and optimally chosen messages—the sender faces a canonical Bayesian persuasion problem with value of posteriors $w_{\beta}$. Since $w_{\beta}\geq v_{-}$, we have that $(\operatorname*{cav}w_{\beta})(\overline{p})\geq(\operatorname*{cav}v_{-})(\overline{p})$.
Observe next $v_{-}=w_{\beta}$ on $\operatorname*{supp}\pi_{S}$, from Theorem 1. It follows that, in an auxiliary Bayesian persuasion problem in which the sender faces value of posteriors $v_{-}$, the sender can secure a payoff of at least $(\operatorname*{cav}w_{\beta})(\overline{p})$ by choosing signal $\pi$. It follows that $(\operatorname*{cav}w_{\beta})(\overline{p})\leq(\operatorname*{cav}v_{-})(\overline{p})$.
3.3 What if the sender cannot prove news better than the prior?
In this case, the stark conflict of interest between the two parties eliminates the possibility of valuable persuasion for the sender.
Proposition 1.
@@ 3.3 What if the sender cannot prove news better than the prior?
If the sender cannot prove news better than the prior, there is a sender-preferred equilibrium in which she acquires no information and obtains an expected payoff of $v(\overline{p})$.
That is, in equilibrium the sender can never obtain an expected payoff above the one she would achieve with no information acquisition (and transmission). Furthermore, this expected payoff can be attained in an equilibrium in which the sender indeed acquires (and, therefore, transmits) no information. [^27] There might be equilibria in which she obtains an expected payoff strictly below $v(\overline{p})$, however. For example, modify the setting of Section 3.1 by increasing the prior from $2/5$ to $9/20$, so that the sender can no longer prove news better than the prior. Observe that there is an equilibrium in which the receiver is maximally skeptical following any message and the sender obtains an expected payoff strictly below $v(\overline{p})$. An immediate consequence is that in ‘cheap talk’ environments (Example 3) persuasion has no value for the sender.
Proof.
Such an equilibrium is constructed in Lemma 3, Appendix A. To show that it is sender-preferred, we will show that there is no equilibrium in which the sender obtains an expected payoff strictly larger than $v(\overline{p})$. Suppose that $(\pi,\mu,\beta)$ is such an equilibrium. Let $w_{\beta}(s)\equiv\max_{m\in M(s)}v(\beta(m))$ for $s\in[0,1]$. If the sender obtains a expected payoff strictly larger than $v(\overline{p})$, it must be that $\max_{s\in\operatorname*{supp}\pi_{S}}w_{\beta}(s)>v(\overline{p})$, meaning that, in equilibrium, the receiver must hold a posterior belief strictly above $\overline{p}$ with positive probability, since $v$ is non-decreasing. Let $S^{\prime\prime}\equiv\operatorname*{arg\,max}_{s\in\operatorname*{supp}\pi_{S}}w_{\beta}(s)$ denote the set of on-path sender types obtaining the largest on-path payoff and let $s^{\prime\prime}\equiv\max S^{\prime\prime}$ denote its largest element. Observe also that, since the receiver follows Bayes’ rule on-path, we have that
$w_{\beta}(s)\leq v(\overline{p})\quad\text{for some}\quad s\in\operatorname*{supp}\pi_{S},$ (2)
as necessarily the receiver must hold a posterior belief strictly below the prior with positive probability if she is to hold a posterior belief strictly above the prior with positive probability.
Observe next that $v(s^{\prime\prime})>v(\overline{p})$. If otherwise $v(s^{\prime\prime})\leq v(\overline{p})$, we would have that $v(s)\leq v(\overline{p})$ for all $s\in S^{\prime\prime}$, since $v$ is non-decreasing and $s^{\prime\prime}=\max S^{\prime\prime}$; however, by construction, $w_{\beta}(s)>v(\overline{p})$ for all $s\in S^{\prime\prime}$, contradicting that the receiver updates using Bayes’ rule on-path.
$v(s^{\prime\prime})>v(\overline{p})$ in turn implies that $s^{\prime\prime}>\overline{p}$ (since $v$ is non-decreasing) and that $s^{\prime\prime}$ is not lowest-consistent (since the sender cannot prove news better than the prior). This latter fact means that for any fixed $m\in\operatorname*{supp}\mu(\cdot|\pi,s^{\prime\prime})$ there exists an $s^{\prime}<s^{\prime\prime}$ such that $m\in M(s^{\prime})$.
Suppose first that $s^{\prime}<\overline{p}$ (alternatively, $s^{\prime}=\overline{p}$), then the sender can obtain a payoff of $w_{\beta}(s^{\prime\prime})>v(\overline{p})$ with probability $1$ by deviating to a signal $\pi^{\prime}$ such that $\operatorname*{supp}\pi_{S}^{\prime}=\{s^{\prime},s^{\prime\prime}\}$ (alternatively, $\operatorname*{supp}\pi_{S}^{\prime}=\{\overline{p}\}$). Since (2) must hold, this is a profitable deviation, so we cannot be at an equilibrium.
Suppose next that $s^{\prime}\in(\overline{p},s^{\prime\prime})$ and consider the following deviating signal $\pi^{\prime}$. Let $\pi_{S}^{\prime}(s^{\prime\prime})=0$ and define $k\equiv(s^{\prime}-\overline{p})+(s^{\prime\prime}-s^{\prime})\pi_{S}(s^{\prime\prime})$. Let
$\pi_{S}^{\prime}(s^{\prime})\equiv\frac{s^{\prime\prime}-\overline{p}}{k}\pi_{S}(s^{\prime\prime})+\frac{s^{\prime}-\overline{p}}{k}\pi_{S}(s^{\prime})>\pi_{S}(s^{\prime\prime})+\pi_{S}(s^{\prime})$
and
$\pi_{S}^{\prime}(s)\equiv\frac{s^{\prime}-\overline{p}}{k}\pi_{S}(s)<\pi_{S}^{\prime}(s)$
for $s\in\operatorname*{supp}\pi_{S}\setminus\{s^{\prime},s^{\prime\prime}\}$. By construction, $\pi_{S}^{\prime}$ is the marginal of a signal. Since $\pi_{S}^{\prime}$ strictly increases the probability assigned to the set $\{s^{\prime},s^{\prime\prime}\}$ (attaining the highest on-path payoff $w_{\beta}(s^{\prime\prime})$, by construction) while lowering the probability on the rest of the support (which leads to an average payoff strictly below $w_{\beta}(s^{\prime\prime})$, since (2) holds), it constitutes a profitable deviation.
@@ 4 Verifiability comparative statics
How does the sender’s equilibrium value change as the extent to which she can misreport her acquired information varies? Which shifts towards ‘more verifiability’ are desirable from the sender’s perspective?
@@
I then show that the result is tight: if the sender cannot prove news better than the prior, increases in verifiability may—perhaps surprisingly—hurt her in equilibrium. I conclude the section by characterizing verifiability structures that are optimal for the sender.
@@ 4.1 More verifiability
As mentioned in Example 4, we can interpret interval-partition verifiability structures as scoring protocols: the sender can prove (perhaps via a third-party certifier) that her type lies in a certain interval which corresponds to a score. It turns out that, if the sender chooses a certifier with a finer grid of scores (or if regulation imposes such a refinement), we have a $\succeq^{\mathrm{lc}}$-increase in the verifiability structure. In this context, the comparative statics question can be cast as: when will the sender benefit, ex ante, for a finer certification technology?
Formally, for some given left-closed partition $\mathcal{P}$ of $[0,1]$ (that is: every $P\in\mathcal{P}$ is left-closed) let $M_{\mathcal{P}}$ denote a partitional verifiability structure as defined in Example 4. If $\mathcal{P}^{\prime\prime}$ is a refinement of $\mathcal{P}^{\prime}$ then $M_{\mathcal{P}^{\prime\prime}}\succeq^{\mathrm{lc}}M_{\mathcal{P}^{\prime}}$. For interval partitions, also the converse holds: if $\mathcal{P}^{\prime\prime}$ is a not a refinement of $\mathcal{P}^{\prime}$ then $M_{\mathcal{P}^{\prime\prime}}\not\succeq^{\mathrm{lc}}M_{\mathcal{P}^{\prime}}$. $\diamond$
@@ 4.2 Comparative statics result
Let $W_{M}(v,\overline{p})$ denote the equilibrium payoff set for admissible parameters $(v,\overline{p},M)$. Write $w_{M}(v,\overline{p})$ when the equilibrium payoff set is a singleton, as in the case when the sender can prove news better than the prior (Corollary 2).
Theorem 2(a) (More verifiability: sufficiency).
Let $(v,\overline{p},M^{\prime})$ be such that the sender can prove news better than the prior. Then
$M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}\quad\Rightarrow\quad w_{M^{\prime\prime}}(v,\overline{p})\geq w_{M^{\prime}}(v,\overline{p}).$
Theorem 2(b) (More verifiability: necessity).
If $M^{\prime\prime}\not\succeq^{\mathrm{lc}}M^{\prime}$ then there exists a triple $(v,\overline{p},M^{\prime})$ such that the sender can prove news better than the prior and
$\sup W_{M^{\prime\prime}}(v,\overline{p})<w_{M^{\prime}}(v,\overline{p}).$
$\succeq^{\mathrm{lc}}$ is therefore the appropriate pre-order to compare verifiability structures from the sender’s perspective, provided that she can prove news better than the prior: changes in the verifiability structure that do not increase (in the set-inclusion sense) which types are lowest-consistent may hurt the sender. The result can be directly applied to compare the verifiability structures introduced in the context of the examples.
@@
If the sender can prove news better than the prior, ‘adding a message’ to any verifiability structure (as defined in Example 7) is (weakly) beneficial for her. That is also the case for refining partitional structures (as introduced in Example 4 and discussed further in Example 6). For partitional structures, the converse applies to interval partitions: if an interval partition is not a refinement of another, the sender will obtain a strictly better equilibrium outcome in the latter, for some admissible $v$ and $\overline{p}$.
Theorem 2(a)-2(b) also characterizes which changes in the sender’s commitment to reveal acquired information are valuable for her. In this interpretation, following signal realization $s$, the sender can commit to using a message from the set $M(s)$. Changes in $M$ can therefore be viewed as changes in the sender’s commitment power. In this context, the theorem characterizes which shifts in the sender’s ability to commit are beneficial for her.
Proof.
Start with 2(a). We use a replication argument which exploits Corollary 1. Consider first the triple $(v,\overline{p},M^{\prime})$ and any equilibrium signal $\pi^{\prime}$. We will show that, in any equilibrium with parameters $(v,\overline{p},M^{\prime\prime})$, the sender’s expected payoff from $\pi^{\prime}$ is at least as large as her (equilibrium) expected payoff from $\pi^{\prime}$ with parameters $(v,\overline{p},M^{\prime})$. This implies that the equilibrium expected payoff with $(v,\overline{p},M^{\prime\prime})$ must also exceed the equilibrium expected payoff under $(v,\overline{p},M^{\prime})$. To see this, observe that since the sender can prove news better the prior, in any equilibrium with $(v,\overline{p},M^{\prime})$ Corollary 1 implies that at any $s\in\operatorname*{supp}\pi_{S}$ the sender obtains a continuation payoff of $v(s)$. Since $s$ is lowest-consistent under $M^{\prime}$ and $L_{M^{\prime\prime}}\supseteq L_{M^{\prime}}$, $s$ is also lowest-consistent under $M^{\prime\prime}$. This in turn implies that, for any equilibrium $(\pi^{\prime\prime},\mu^{\prime\prime},\beta^{\prime\prime})$ when parameters are $(v,\overline{p},M^{\prime\prime})$, $\max_{m\in M^{\prime\prime}(s)}v(\beta^{\prime\prime}(m))\geq v(s)$, as each $s\in\operatorname*{supp}\pi_{S}$ is lowest-consistent by Theorem 1. Hence the sender’s expected payoff from $\pi^{\prime}$ with parameters $(v,\overline{p},M^{\prime\prime})$ is at least as large as her (equilibrium) expected payoff from $\pi^{\prime}$ with parameters $(v,\overline{p},M^{\prime})$, as required.
Now to the converse 2(b). Observe first that $M^{\prime\prime}\not\succeq^{\mathrm{lc}}M^{\prime}$ implies that $L_{M^{\prime}}\neq\{0\}$, as $0\in L_{M}$ for any $M$. Since $L_{M^{\prime}}\neq\{0\}$, there exists some $s^{*}\in L_{M^{\prime}}$, $s^{*}>0$ such that $s^{*}\notin L_{M^{\prime\prime}}$. Take $v(s)=1(s\geq s^{*})$ and $\overline{p}=s^{*}/2$. Let $v_{-}^{\prime}$ and $v_{-}^{\prime\prime}$ denote the value of posteriors with maximal receiver skepticism as defined in equation (1) when the verifiability structures are $M^{\prime}$ and $M^{\prime\prime}$, respectively. As the sender can prove news better than the prior with parameters $(v,\overline{p},M^{\prime})$, she obtains her full-commitment payoff $(\operatorname*{cav}v)(\overline{p})$ in equilibrium (observing that $(\operatorname*{cav}v)(\overline{p})=(\operatorname*{cav}v_{-}^{\prime})(\overline{p})$ and applying Corollary 2). This expected payoff is not attainable in equilibrium under $M^{\prime\prime}$: if the sender cannot prove news better than the prior under $M^{\prime\prime}$, this follows from Proposition 1; if the sender can prove news better than the prior under $M^{\prime\prime}$; this follows from observing that $(\operatorname*{cav}v)(\overline{p})>(\operatorname*{cav}v_{-}^{\prime\prime})(\overline{p})$ and applying Corollary 2.
@@
An alternative (natural) notion of ‘more verifiability’ captures the idea that, for all sender types, separation possibilities increase. Formally, for a fixed verifiability structure $M$ and message $m\in\mathcal{M}$, let $S_{m}^{c}\equiv[0,1]\setminus M^{-1}(m)$ denote the complement of $M^{-1}(m)$ in $[0,1]$ i.e. the set of types who cannot send message $m$. For a given $s\in[0,1]$, let
@@
It can be easily verified that $M^{\prime\prime}\succeq M^{\prime}$Rightarrow $M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}$ while $M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}$$\not\Rightarrow$ $M^{\prime\prime}\succeq M^{\prime}$. Shifts to a $\succeq$-higher structure are therefore sufficient to weakly improve the sender’s equilibrium payoff (Theorem 2(a)) (provided she can prove news better than the prior) but they are not necessary (as highlighted in Theorem 2(b)) because of the monotonicity of $v$.

### v6

@@ 3.2 Proving news better than the prior
Being able to prove news better than the prior means that the sender can—by acquiring information in a way that makes sending message $m_{h}$ possible—induce a belief in the receiver that she strictly prefers over the prior. Importantly, this is possible irrespective of the receiver’s belief (i.e., even if she is maximally skeptical, assigning belief $\min M^{-1}(m_{h})$ to message $m_{h}$).
@@
PNBP can hold only if, in the terminology of Kamenica and Gentzkow (2011), there is information the sender would share. That is, only if the sender is not already obtaining her highest possible payoff at the prior. Notice that in full-verifiability settings (Examples 1 and 2) this condition is also sufficient for PNBP. In contrast, in cheap talk environments (Example 3) the sender cannot prove news better than the prior for any admissible $v$ and $\overline{p}$ (so even if there is information she would share).
Theorem 2 (PNBP, unraveling and persuasion).
@@
If the sender can prove news better than the prior, every equilibrium is unraveling. If instead the sender cannot prove news better than the prior, there exists a sender-preferred unpersuasive equilibrium in which she acquires no information.
The proof is in Appendix C. Theorem 2 implies that, if the sender can prove news better than the prior, her equilibrium expected payoff is unique, because of Corollary 1.191919PNBP does not ensure that the sender’s expected payoff in equilibrium is at least $v(\overline{p})$. For instance, in the example presented in Section 4.3 verifiability structure $M^{\prime\prime}$ leads to a unique equilibrium which is unraveling and such that the sender’s expected payoff is strictly below $v(\overline{p})$. In contrast, if the sender cannot prove news better than the prior, any equilibrium which is not unpersuasive—i.e., which is unraveling, in light of Theorem 1—leads to an expected sender payoff not higher than $v(\overline{p})$. [^20] It is straightforward to construct examples of such unraveling equilibria. For example, modify the example in Section 1.1 by increasing the prior from $1/4$ to $9/20$, so that PNBP no longer holds. In this case, there is an unraveling equilibrium in which the receiver is maximally skeptical following every message and the sender obtains an expected payoff strictly below $v(\overline{p})$.
When PNBP fails, therefore, the sender’s bias in favor of state $1$ eliminates the possibility of valuable persuasion. That is, in equilibrium the sender can never obtain an expected payoff above the one she would achieve with no information acquisition (and transmission). An immediate consequence is that in ‘cheap talk’ environments (Example 3) covert learning has no value for the sender.
Why is PNBP sufficient for unraveling? In light of Theorem 1, it is enough to argue that if PNBP holds and an equilibrium is unpersuasive, then it must also be unraveling. The intuition is as follows. PNBP guarantees the existence of a message conveying ‘good news’—that is, a message that yields the sender a payoff strictly above $v(\overline{p})$, even under maximal receiver skepticism. In an unpersuasive equilibrium, the sender must obtain $v(\overline{p})$ with probability one, so she must be deterred from deviating to a strategy in which she acquires information, sends the ‘good news’ message when she indeed discovers good news, and otherwise falls back on an on-path message that delivers $v(\overline{p})$.
Such a deviation can be ruled out only if, in equilibrium, the sender learns nothing, and the prior itself is lowest-consistent with some message $m$ the sender uses. In this case, $m$ can no longer serve as a fallback in the deviation just described. But then, the unpersuasive equilibrium is also an unraveling one.
To conclude this section and illustrate its results, it is informative to consider their implications in a setting with Dye verifiability, as introduced in Example 5.
Example 5 (Dye verifiability, continued).
Assume that $v(p)=p$, so that the sender obtains a payoff equal to the expected value of the state. This specialization of the model brings it closest to the set-up studied in DeMarzo et al. (2019). Assume, to rule out trivial cases, that $\overline{p}\in(0,1)$. My results directly imply the following: (i) the sender’s equilibrium expected payoff is equal to $v(\overline{p})$, (ii) the receiver’s belief following the non-disclosure message $m_{\emptyset}$ is equal to 0 in all equilibria.
Point (i) holds because an equilibrium exists [^21] Lemma 1 cannot be invoked for equilibrium existence, since the sufficient condition it relies on fails in this example. It is immediate that an equilibrium exists, however. For example, this is an equilibrium: the sender perfectly learns the value of the state (so $\pi_{S}(1)=\overline{p}$ and $\pi_{S}(0)=1-\overline{p}$) and sends message $s$ when her type is $s\neq\overline{p}$; the receiver’s belief is $\beta(m_{\emptyset})=0$ and $\beta(s)=s$ for $s\in[0,1]\setminus\{\overline{p}\}$. and all equilibria are unraveling (because PNBP holds, so Theorem 2 applies) so they all lead to the same expected payoff for the sender, which is equal to $(\operatorname*{cav}\underline{v}_{M})(\overline{p})=v(\overline{p})$, by Corollary 1.
To see why (ii) holds, suppose otherwise that the receiver’s belief following the non-disclosure message $m_{\emptyset}$ were equal to some $\beta(m_{\emptyset})>0$ in equilibrium, implying that the sender’s payoff from sending message $m_{\emptyset}$ would strictly exceed $v(0)$. The sender could then deviate to acquiring full information, revealing what she learns when $\omega=1$ (by sending message 11—recall that $M(1)=\{1,m_{\emptyset}\}$) and withholding her findings when $\omega=0$ (by sending message $m_{\emptyset}$—recall that $M(0)=\{0,m_{\emptyset}\}$). This would lead to an expected payoff equal to
$\overline{p}v(1)+(1-\overline{p})v(\beta(m_{\emptyset}))>\overline{p}v(1)+(1-\overline{p})v(0)=v(\overline{p}),$
meaning that there cannot be an equilibrium in which the sender’s expected payoff is equal to $v(\overline{p})$, contradicting (i).
These two observations correspond to the ‘minimum principle’ in DeMarzo et al. (2019) (their Proposition 1). The most substantial difference between this specialization of my model and theirs is that they consider a setting in which the sender is constrained in her information acquisition choice in the sense that all available signals put positive probability on the unverifiable signal realization. Therefore, in their model the non-disclosure posterior belief may strictly exceed zero. They show that this belief must be minimal, in equilibrium. In contrast, my sender is able to choose any signal, so minimality of the non-disclosure posterior belief means that it must be equal to zero.  $\diamond$
@@ 4 Verifiability comparative statics
How does the sender’s equilibrium payoff change as the extent to which she can misreport her acquired information varies? Which shifts towards ‘more verifiability’ are desirable from the sender’s perspective?
@@
I then show that the result is tight: if the sender cannot prove news better than the prior, meaning that unpersuasive equilibria exist, increases in verifiability may—perhaps surprisingly—hurt her in equilibrium. I conclude the section by characterizing verifiability structures that are optimal for the sender.
@@ 4.1 More verifiability
As mentioned in Example 4, we can interpret interval-partition verifiability structures as scoring protocols: the sender can prove (perhaps via a third-party certifier) that her type lies in a certain interval, which corresponds to a score. It turns out that, if the sender chooses a certifier with a finer grid of scores (or if regulation imposes such a refinement), we have a $\succeq^{\mathrm{lc}}$-increase in the verifiability structure. In this context, the comparative statics question can be cast as: when will the sender benefit, ex ante, from a finer certification technology?
Formally, for some given left-closed partition $\mathcal{P}$ of $[0,1]$ (that is: every $P\in\mathcal{P}$ is left-closed) let $M_{\mathcal{P}}$ denote a partitional verifiability structure as defined in Example 4. If $\mathcal{P}^{\prime\prime}$ is a refinement of $\mathcal{P}^{\prime}$ then $M_{\mathcal{P}^{\prime\prime}}\succeq^{\mathrm{lc}}M_{\mathcal{P}^{\prime}}$. For interval partitions, also the converse holds: if $\mathcal{P}^{\prime\prime}$ is not a refinement of $\mathcal{P}^{\prime}$ then $M_{\mathcal{P}^{\prime\prime}}\not\succeq^{\mathrm{lc}}M_{\mathcal{P}^{\prime}}$.  $\diamond$
@@ 4.2 Comparative statics result
Let $\underline{V}_{M}^{*}(v,\overline{p})$ and $\overline{V}_{M}^{*}(v,\overline{p})$ denote the lowest and highest equilibrium expected payoffs attainable for the sender at triple $(v,\overline{p},M)$, respectively. To ensure that these values are well-defined, for the rest of this section we will restrict attention to verifiability structures satisfying the continuity requirement stated in the existence lemma (Lemma 1).
Recall that Theorem 2 and Corollary 1 together imply that $\overline{V}_{M}^{*}(v,\overline{p})=\underline{V}_{M}^{*}(v,\overline{p})=(\operatorname*{cav}\underline{v}_{M})(\overline{p})$ if PNBP holds and $\overline{V}_{M}^{*}(v,\overline{p})=v(\overline{p})$ if it does not.
Theorem 3 (More verifiability).
Consider verifiability structures $M^{\prime}$ and $M^{\prime\prime}$. Then
$M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}\quad\Leftrightarrow\quad\overline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})\geq\overline{V}_{M^{\prime}}^{*}(v,\overline{p})$
for every admissible $v$ and $\overline{p}$ such that at $(v,\overline{p},M^{\prime})$ the sender can prove news better than the prior.
The proof is in Appendix D. The result can be directly applied to compare the verifiability structures introduced in the context of the examples.
@@
If the sender can prove news better than the prior, ‘adding a message’ to any verifiability structure (as defined in Example 8) is beneficial for her. That is also the case for refining partitional structures (as introduced in Example 4 and discussed further in Example 7). For partitional structures, the converse applies to interval partitions (see, again, Example 4 for a definition): if an interval partition is not a refinement of another, the sender will obtain a strictly better equilibrium outcome in the latter, for some admissible $v$ and $\overline{p}$.
Theorem 3 also characterizes which changes in the sender’s commitment to reveal acquired information are valuable for her. In this interpretation, following signal realization $s$, the sender can commit to using a message from the set $M(s)$. Therefore, $\succeq^{\mathrm{lc}}$-increases in $M$ are exactly the changes in the sender’s commitment power which are beneficial for her.
Theorem 3 follows straightforwardly from the following lemma of independent interest (also proved in Appendix D).
Lemma 2.
Consider verifiability structures $M^{\prime}$ and $M^{\prime\prime}$. Then
$M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}\quad\Leftrightarrow\quad\underline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})\geq\underline{V}_{M^{\prime}}^{*}(v,\overline{p})$
for every admissible $v$ and $\overline{p}$ such that at $(v,\overline{p},M^{\prime})$ an unraveling equilibrium exists.
Therefore, whenever an unraveling equilibrium exists (regardless of whether PNBP holds or not) a $\succeq^{\mathrm{lc}}$-increase in verifiability cannot lower the sender-worst equilibrium payoff. Conversely, changes in the verifiability structure that do not increase (in the set-inclusion sense) which types are lowest-consistent may do so.
@@
An alternative (natural) notion of ‘more verifiability’ captures the idea that, for all sender types, separation possibilities increase. Formally, for a fixed verifiability structure $M$ and message $m\in\mathcal{M}$, let $S_{m}^{c}\equiv[0,1]\setminus M^{-1}(m)$ denote the complement of $M^{-1}(m)$ in $[0,1]$. That is, $S_{m}^{c}$ is the set of types that cannot send message $m$. For a given $s\in[0,1]$, let
@@
It can be easily verified that $M^{\prime\prime}\succeq M^{\prime}$ $\Rightarrow$ $M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}$ while $^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime} ⇏\not\Rightarrow M′′M$$^{\prime\prime}\succeq M^{\prime}$. Shifts to a $\succeq$-higher structure are therefore sufficient to weakly improve the sender’s equilibrium payoff (provided she can prove news better than the prior) but they are not necessary (as highlighted in Theorem 3) because of the monotonicity of $v$.